# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ 3.1. Restriction measures and CLEs
On the other hand, the Conformal Loop Ensembles (CLE) are another one-parameter family of conformally invariant measures, this time on collections of non-nested disjoint simple loops in a domain $D$. There are various ways to construct them (see [SW12]), but we will see them as the collection of the outermost boundaries of clusters in a Brownian loop-soup (i.e. a Poisson point process with respect to the Brownian loop measure) sample with intensity $\mathbf{c}/2=\mathbf{c}(\kappa)/2$, $0<\mathbf{c}\leq 1$. We stress that the intensity parameter in [SW12, Law09] is missing this factor of $1/2$: see also [Lup18] for a more thorough explanation. Note that $1/2$ is the critical identity, above which there is a.s. one cluster in the Brownian loop soup and its boundary coincides with $\partial D$, and that the range $8/3<\kappa\leq 4$ exactly corresponds to $0<\mathbf{c}\leq 1$.
@@
The CLE also satisfies a form of conformal restriction: for $D^{\prime}\subset D$, consider the set $\Gamma^{\prime}$ obtained from a realization $\Gamma$ of the CLE in $D$ consisting of all the loops that intersect $D^{\prime}$, and let $D_{\Gamma}^{\prime}$ be the random set given by the union of $D^{\prime}$ with the filling of $\Gamma$. Then, conditionally on $D_{\Gamma}^{\prime}$, the law of the loops of the CLE that do not intersect $D_{\Gamma}^{\prime}$ is that of a CLE in $D\setminus D_{\Gamma}^{\prime}$.
3.2. The non-crossing $\mathrm{SLE}_{\kappa}$
@@
We work with the notation of the previous section, only redefining the parameters $\mathbf{b}$ to belong to {±$\left\{\pm 1\right\}^{n}$ (the purpose of the factor $\lambda$ was just to make a notational connection with the boundary conditions of the GFFs, but all that matters is the sign of $\mathbf{b}$). We will also consider for $D$ an admissible domain and $\mathbf{b}\in\left\{\pm 1\right\}^{n}$ the harmonic function $u_{D}$ given by the boundary values $-1$ on $\mathcal{B}_{\mathbf{b}}^{-}$, $+1$ on $\mathcal{B}_{\mathbf{b}}^{+}$, so that $u_{\mathbf{b}}=g_{\mathbf{b}}/\lambda$ from the previous section.
@@
Consider $K$ a sample of the restriction measure $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$, $\Gamma$ a sample of an independent $\operatorname{CLE}_{\kappa}^{D}$, and $C$ the union of $K$ with all the loops of $\Gamma$ that it intersects. We let $\mathbb{P}_{D}^{\mathbf{b}}$ be the law of the right boundary of the component of $C$ that contains the arc $(yx)$, i.e. the boundary of its filling minus $(yx)$. This is a simple curve in $D$ (see [WW13] for an argument in the simply connected case: the general case is identical). In other words, $\mathbb{P}_{D}^{\mathbf{b}}=F_{\ast}(\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}})$, where $F$ identifies the path according to the construction above given a realization of the $\operatorname{CLE}$ and of the restriction measure.
Theorem 19.
@@
Let $D\subset\mathbb{D}$ be admissible and $8/3<\kappa\leq 4$. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(28) $\nu_{D}(\eta):=\sum_{\mathbf{b}\in\left\{\pm 1\right\}^{n}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{\mathbf{b}}(\eta)\cdot Z_{D,\mathbf{b}},$
@@
(29) $Z_{D,\mathbf{b}}=\exp(-\frac{\pi h}{4}\norm{u_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2}),$
@@
Note that $\pi h_{4}/4=\pi/16=\lambda^{2}/2$ so that the expression above is compatible with Equation (10).
@@
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\left\{\pm 1\right\}^{n}}\exp(-\frac{\pi h_{\kappa}}{4}\norm{u_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2})\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
@@
The proof of the Theorem will again follow Lawler’s characterization of the $\mathrm{SLE}$ measures. In fact, it will be sufficient to show that for $K\subset D$ a compact at positive distance from $\mathcal{B}_{l}$ such that $\mathcal{B}_{r}\cup K$,
(30) $\frac{\mathrm{d}\nu_{D\setminus K}(x,y)}{\mathrm{d}\nu_{D}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D}\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,K)).$
Indeed, this would characterize $\nu_{D}$ seen as a law on connected hulls containing $(yx)$, i.e. right-sided restriction measures from $x$ to $y$: since the $\mathrm{SLE}_{\kappa}$ measures should have the same Radon-Nikodym derivative, they must correspond to the same right-sided restriction measure, and the fact that their boundary are given by a simple curve is enough to assert that they also coincide as measures on simple curves from $x$ to $y$.
@@
Let $D^{\prime}:=D\setminus K$ be as above. Up to relabeling the boundary components $(\mathcal{B}_{i})_{1\leq i\leq n}$, we may assume that $(\mathcal{B}_{i})_{1\leq i\leq k}$ are the inner boundary components of $D^{\prime}$. We have to verify that $\nu_{D}$ satisfies Equation (30), i.e. that
@@
Because on $\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}$, $\mathbb{P}_{D}^{\mathbf{b}}$ (resp. $\mathbb{P}_{D^{\prime}}^{\mathbf{b}_{k}}$) is a deterministic function of a realization of $\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ (resp. of $\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$), it suffices to compute [^4] We are abusing notation: the equation should be understood as $\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(F(\Gamma,K))=\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}}(F(\cdot,K))|_{\cdot=\Gamma}\frac{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(F(\Gamma,\cdot))|_{\cdot=K}$
@@
On the one hand, the sample of the Brownian loop soup in $D$ of intensity $\mathbf{c}/2$ from which $\operatorname{CLE}_{\kappa}^{D}$ is constructed consists of loops in $D^{\prime}$ and loops intersecting $D^{\prime}$. The former corresponds to a sample of a Brownian loop soup in $D^{\prime}$ and the outermost boundaries of its clusters yield a realization of $\operatorname{CLE}_{\kappa}^{D^{\prime}}$. One can therefore construct a curve $\tilde{\eta}^{\prime}$ from $\operatorname{CLE}_{\kappa}^{D^{\prime}}$, and then explore the boundary of the loops in the remainder of the loop soup in $D$ to obtain $\tilde{\eta}$. Both curves are then such that either $\tilde{\eta}^{\prime}=\tilde{\eta}$, or $\tilde{\eta}\not\subset D^{\prime}$: and the former happens if and only if the loop soup does not contain loops intersecting both $\eta$ and $D\setminus D^{\prime}$. Since the number of these loops is a Poisson random variable with parameter $\mathbf{c}\cdot m_{D}(\eta,D\setminus D^{\prime})/2$, we obtain that
@@
On the other hand, $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ can be obtained by first sampling $R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$, then an independent Poisson point process of Brownian excursions intersecting $D\setminus D^{\prime}$, of intensity $h=h_{\kappa}$. Again here, the curve $\tilde{\eta}^{\prime}$ obtained from $R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$ and the curve $\eta$ obtained from $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ are such that $\tilde{\eta}=\tilde{\eta}^{\prime}$ or $\tilde{\eta}\not\subset D^{\prime}$, the former happening if and only if the realisation of the PPP does not contain excursions intersecting $\mathcal{B}_{r}^{\prime}$ (note that these excursions necessarily hit $\eta$), i.e. if and only if no excursion in the realization of $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ hit $D\setminus D^{\prime}$. Once again we find that
$\frac{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}(\eta)=\mathbb{P}\left[\not\exists\gamma\in R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}:\gamma\cap D\setminus D^{\prime}\neq\emptyset\right]^{-1}=\exp(\frac{\pi h}{4}(\norm{g_{D^{\prime},\mathbf{b}_{k}}}_{\nabla(D^{\prime})}^{2}-\norm{g_{D,\mathbf{b}}}_{\nabla(D)}^{2})),$
using Lemma 18. The right-hand term of this equality cancels out with $Z_{D^{\prime},\mathbf{b}_{k}}/Z_{D,\mathbf{b}}$ and concludes the proof. ∎
@@ Appendix A Existence of level-lines for compactified GFF on crossing admissible domains
Here we explain in more details the existence of level-lines for compactified GFF, in the crossing case. Let $(D,x,y)$ be a crossing admissible domain: suppose first that $D$ is one-connected, and by conformal invariance consider $D=\mathbb{A}_{p}$. Let $\mathbf{b}=b_{\text{in}}\in\mathbb{Z}$ and $g_{b_{\text{in}}}$ as above, the multivalued harmonic function with locally constant boundary conditions with a jump of $+2\lambda$ across $x$ and $y$, and a change of fixed branch of $2\lambda b_{\text{in}}$ from the outer to the inner boundary of $D$. The existence of a level-line for a (compactified) GFF $\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}}$ is then provided by the following result:
Proposition 21 ([IK13, Proposition 10]).

### v2

@@ 3.1.2. Conformal Loop Ensembles
The Conformal Loop Ensembles (CLE) are another one-parameter family of conformally invariant measures, this time on collections of non-nested disjoint simple loops in a domain $D$ for the parameter range $8/3<\kappa\leq 4$. There also admit various constructions (see [37]), but we will see them here as the collection of the outermost boundaries of clusters in a Brownian loop-soup, i.e. a Poisson point process with respect to the Brownian loop measure, with intensity $\alpha=\mathbf{c}(\kappa)/2$, $0<\alpha\leq 1/2$. We stress that the intensity parameter in [37, 24] and the central charge $\mathbf{c}$ differ by a factor of $1/2$: see also [29] for a more thorough explanation.
@@
The CLE also satisfies a form of conformal restriction: for $D^{\prime}\subset D$, consider the set $\Gamma^{\prime}$ obtained from a realisation $\Gamma$ of the CLE in $D$ consisting of all the loops that intersect $D^{\prime}$, and let $D_{\Gamma}^{\prime}$ be the random set given by the union of $D^{\prime}$ with the filling of $\Gamma^{\prime}$. Then, conditionally on $D_{\Gamma}^{\prime}$, the law of the loops of the CLE that do not intersect $D_{\Gamma}^{\prime}$ is that of a CLE in $D\setminus D_{\Gamma}^{\prime}$.
3.2. Constructing the non-crossing $\mathrm{SLE}_{\kappa}$
@@
We work with the notation of the previous section, only redefining the parameters $\mathbf{b}$ to belong to $\left\{\pm 1\right\}^{n}$ (the purpose of the factor $\lambda$ was just to make a notational connection with the boundary conditions of the GFFs, but all that matters is the sign of $\mathbf{b}$). We will also consider for $D$ an admissible domain and $\mathbf{b}\in\left\{\pm 1\right\}^{n}$ the harmonic function $u_{\mathbf{b}}$ given by the boundary values 1-1 on $\mathcal{B}_{\mathbf{b}}^{-}$, +1+1 on $\mathcal{B}_{\mathbf{b}}^{+}$, so that $u_{\mathbf{b}}=g_{\mathbf{b}}/\lambda$ from the previous section.
@@
Consider $R$ a sample of $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$, $\Gamma$ a sample of an independent $\operatorname{CLE}_{\kappa}^{D}$, and $C$ the union of $R$ with all the loops of $\Gamma$ that it intersects. We let $\mathbb{P}_{D}^{\mathbf{b}}$ be the law of the right boundary of the component of $C$ that contains the arc $(yx)$, i.e. the boundary of its filling minus $(yx)$. This is a simple curve in $D$ (see [44], just before Section 3, for an argument in the simply connected case: the general case is identical). In other words, $\mathbb{P}_{D}^{\mathbf{b}}=F_{\ast}(\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}})$, where $F$ identifies the path according to the construction above given a realisation of the $\operatorname{CLE}$ and of the PPP of Brownian excursions.
Theorem 3.2.
@@
Let $(D,x,y)$ be a non-crossing admissible domain and $8/3<\kappa\leq 4$. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(33) $\nu_{D}:=\sum_{\mathbf{b}\in\left\{\pm 1\right\}^{n}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{\mathbf{b}}\cdot Z_{D,\mathbf{b}},$
@@
(34) $Z_{D,\mathbf{b}}=\exp(-\frac{\pi h}{4}\norm{u_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2}),$
@@
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\left\{\pm 1\right\}^{n}}\exp(-\frac{\pi h}{4}\norm{u_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
@@
As a minor check we note that $\pi h_{4}/4=\pi/16=\lambda^{2}/2$ and hence the expression above agrees with Equation (14) when $\kappa=4$.
3.2.1. Proof of Theorem 3.2
The proof of the theorem will again just check the conditions from Lawler’s characterisation of the $\mathrm{SLE}$ measures. Among those, it is the restriction property that needs checking. To do this, it is in fact sufficient to show that
(35) $\frac{\mathrm{d}\nu_{D\setminus K}(x,y)}{\mathrm{d}\nu_{D}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D}\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,K))$
for $K\subset\overline{D}$ a hull attached to $\mathcal{B}_{r}$, i.e. such that $K=\overline{K\cap D}$, $\overline{K}\cap\mathcal{B}_{l}=\emptyset$, $D\setminus K$ and $K\cup\partial\mathbb{D}$ connected. Indeed, this would characterise $\nu_{D}$ when seen as a law on connected hulls containing $(yx)$ (the left filling of its samples): since the $\mathrm{SLE}_{\kappa}$ measures should have the same Radon-Nikodym derivative, they must correspond to the same law on connected hulls containing $(yx)$, and the fact that they are simple curve—i.e. they coincide with their right boundary—is enough to assert that $\nu_{D}$ and $\mathrm{SLE}_{\kappa}$ also coincide as measures on simple curves from $x$ to $y$.
@@
Let $D^{\prime}:=D\setminus K$ be as above. Up to relabeling the boundary components $(\mathcal{B}_{j})_{1\leq j\leq n}$ of $D$, we may assume that $(\mathcal{B}_{j})_{1\leq j\leq k}$ are the inner boundary components of $D^{\prime}$. We have to verify that $\nu_{D}$ satisfies Equation (35), i.e. that
@@
Given that on $\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}$, $\mathbb{P}_{D}^{\mathbf{b}}$ (resp. $\mathbb{P}_{D^{\prime}}^{\mathbf{b}_{k}}$) is a deterministic function of a realisation of $\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ (resp. of $\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$), it suffices to compute [^12] We are abusing notation: the equation should be understood as $\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(F(\Gamma,K))=\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}}(F(\cdot,K))|_{\cdot=\Gamma}\frac{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(F(\Gamma,\cdot))|_{\cdot=K}$
@@
On the one hand, the sample of the Brownian loop soup in $D$ of intensity $\mathbf{c}/2$ from which $\operatorname{CLE}_{\kappa}^{D}$ is constructed consists of loops in $D^{\prime}$ and loops intersecting $D^{\prime}$. The former corresponds to a sample of a Brownian loop soup in $D^{\prime}$ and the outermost boundaries of its clusters yield a realisation of $\operatorname{CLE}_{\kappa}^{D^{\prime}}$. One can therefore construct a curve $\tilde{\eta}^{\prime}$ from $\operatorname{CLE}_{\kappa}^{D^{\prime}}$, and then explore the boundary of the loops in the remainder of the loop soup in $D$ to obtain $\tilde{\eta}$. Both curves are then such that either $\tilde{\eta}^{\prime}=\tilde{\eta}$, or $\tilde{\eta}\not\subset D^{\prime}$: and the former happens if and only if the loop soup does not contain loops intersecting both $\tilde{\eta}$ and $D\setminus D^{\prime}$. Since the number of these loops is a Poisson random variable with parameter $\mathbf{c}\cdot m_{D}(\tilde{\eta},D\setminus D^{\prime})/2$, we obtain that
@@
On the other hand, $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ can be obtained by first sampling $R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$, then an independent Poisson point process of Brownian excursions intersecting $D\setminus D^{\prime}$ of intensity $h=h_{\kappa}$. Again here, the curve $\tilde{\eta}^{\prime}$ obtained from $R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$ and the curve $\tilde{\eta}$ obtained from $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ are such that $\tilde{\eta}=\tilde{\eta}^{\prime}$ or $\tilde{\eta}\not\subset D^{\prime}$, the former happening if and only if the realisation of the PPP does not contain excursions intersecting $\mathcal{B}_{r}^{\prime}$ (note that these excursions necessarily hit $\eta$), i.e. if and only if no excursion in the realisation of $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ hit $D\setminus D^{\prime}$. Once again we find that
$\frac{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}(\eta)=\mathbb{P}\left[\not\exists\gamma\in R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}:\gamma\cap D\setminus D^{\prime}\neq\emptyset\right]^{-1}=\exp(\frac{\pi h}{4}(\norm{g_{D^{\prime},\mathbf{b}_{k}}}_{\nabla(D^{\prime}),\mathrm{reg}}^{2}-\norm{g_{D,\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})),$
using Lemma 3.1. The right-hand term of this equality cancels out with $Z_{D^{\prime},\mathbf{b}_{k}}/Z_{D,\mathbf{b}}$ and concludes the proof. ∎
@@ Appendix A Existence of level-lines for compactified GFF on crossing admissible domains
Here we explain in more details the existence of level-lines for the compactified GFF, in the crossing case: the argument is almost identical to the proof of Lemma 15 in [4]. We replicate it here for the sake of completeness.
Proposition A.1.